# Adaptive Monitoring of Stochastic Fire Front Processes via Information-seeking Predictive Control

### v1

@@ I Introduction
The 2025 Southern California wildfires were among the most devastating in the state’s history: the Palisades Fire in Los Angeles and the Eaton Fire in Altadena burned nearly 40,000 acres, destroyed over 16,000 structures, and displaced hundreds of thousands of people [7]. Accurate estimation of wildfire propagation is therefore critical for effective disaster response and informed decision-making. Motivated by this need, this work investigates adaptive monitoring of a stochastic wildfire front using a mobile agent.
The task requires planning the agent’s trajectory over a rolling finite horizon to minimize uncertainty in estimating the fire’s evolution from sensor data. At each step, the agent re-plans based on the current environment, yielding a complex problem that demands an integrated approach to sensing, estimation, and control. Existing methods often address only parts of this problem, either by decoupling sensing, estimation, and control, or by simplifying assumptions [11, 15]. The problem considered in this work relates to active sensing [9], information gathering [1], and sensor management [5]. Sensor management approaches typically address stateless sensors without dynamics, focusing on placement or node selection, and are therefore limited in scenarios where sensor states evolve (e.g., drones equipped with onboard sensors). In such dynamic settings, adaptive control strategies are required. For stochastic dynamic processes, many methods adopt myopic control [3], whereas non-myopic schemes are typically greedy [13] or heuristic with sub-optimality guarantees [6]. Approaches that provide optimality guarantees often assume linear dynamics with Gaussian noise and apply the certainty equivalence principle (CEP), thereby reducing the problem to deterministic optimal control [8]. These assumptions, however, do not hold for the problem addressed in this work.
@@ II-A Problem Objective
Here, the subscript $t+\tau|t$ denotes predicted quantities at time step $t+\tau$ for $\tau\in\{1,\dots,T\}$ within the planning horizon, based on information available at time step $t$. The cost function $\mathcal{C}_{\tau}$ is a bounded, real-valued function that takes as input: a) the predictive density $\mathcal{B}^{-}_{t+\tau|t}$ i.e., $\mathcal{B}^{-}_{t+\tau|t}(X_{t+\tau|t}|Z_{1:t+\tau-1|t})$, of the fire front state at time $t+\tau|t$, and b) the future i.e., predicted, measurement $Z_{t+\tau|t}(u_{t+\tau-1|t})$, which will be received given that the agent executes the control input $u_{t+\tau-1|t}$ and moves to the predicted state $y_{t+\tau|t}$. It then returns a value representing the uncertainty of the fire front state captured in the resulting (pseudo) posterior belief $\mathcal{B}_{t+\tau|t}$. An information-rich measurement set is one that reduces the dispersion in the posterior, which is the agent’s objective over the planning horizon. The expectation is taken with respect to the future measurement set $Z_{t}^{1:T}=\{Z_{t+1|t},\ldots,Z_{t+T|t}\}$. Subsequently, the agent executes the first control input in the sequence, i.e., $u_{t|t}$, transitions to its new state $y_{t+1}$, receives the real measurement $Z_{t+1}$, computes the posterior belief $\mathcal{B}_{t+1}(X_{t+1}|Z_{1:t+1})$, and repeats the process described above for time step $t+1$.
@@ II-B Fire Front Propagation Model
where $SC(i)=\text{x}^{i}_{s}\sin\left(\theta(i)\right)+\text{y}^{i}_{s}\cos\left(\theta(i)\right)$, $CS(i)=\text{x}^{i}_{s}\cos\left(\theta(i)\right)-\text{y}^{i}_{s}\sin\left(\theta(i)\right)$, $C_{1}(i)=\alpha_{3}(i)\sin\left(\theta(i)\right)$, $C_{2}(i)=\alpha_{3}(i)\cos\left(\theta(i)\right)$, and $\text{x}^{i}_{s},\text{y}^{i}_{s}]^{\top}$ are the components of the tangent vector at vertex $i$, providing the local orientation of the fire front at that point.
Environmental conditions, such as fuel type and weather, local to each vertex, affect the forward fire propagation rate and direction. These factors include wind direction and speed, denoted by $\theta$ and $w_{s}$, respectively, as well as the fire spread rate due to fuel type, denoted by $r_{f}$. These parameters are stochastic and can vary throughout the environment. Therefore, each vertex may be affected differently depending on the fire front’s extent and the environmental variability. Specifically, we denote $\theta(i)\in[0,2\pi]$ as the wind direction affecting vertex $i$, the wind speed at the location of vertex $i$ as $w_{s}(i)\in\mathbb{R}^{+}$, and the fire spread rate as $r_{f}(i)\in\mathbb{R}^{+}$. Here, $\theta(i),w_{s}(i),r_{f}(i)$, $\forall i$, are random realizations of the wind direction, wind speed, and fire spread rate at the location of vertex $i$. The parameters $\alpha_{1}(i)$, $\alpha_{2}(i)$, and $\alpha_{3}(i)$ denote the shape parameters governing the elliptical fire growth from vertex $i$, representing respectively the lengths of the semi-minor axis, the semi-major axis, and the distance from the ignition point to the center of the ellipse, defined respectively as: $\alpha_{1}(i)=\frac{r_{f}(i)+\frac{r_{f}(i)}{HB(i)}}{2LB(i)}$, $\alpha_{2}(i)=\frac{r_{f}(i)+\frac{r_{f}(i)}{HB(i)}}{2}$, and $\alpha_{3}(i)=\alpha_{2}(i)-\frac{r_{f}(i)}{HB(i)}$, where $HB(i)$ is the head-to-back ratio accounting for the difference between the fire’s forward (head) and backward (back) spread from the ignition point, while $LB(i)$ is the length-to-breadth ratio which determines the overall elongation of the fire’s elliptical shape. These are defined as: $HB(i)=\frac{LB(i)+(LB(i)^{2}-1)^{0.5}}{LB(i)-(LB(i)^{2}-1)^{0.5}}$, and $LB(i)=0.936\exp(0.2566w_{s}(i))+0.461\exp(-0.1548w_{s}(i))-0.397$. For a more in-depth description these parameters we refer the reader to [4]. Subsequently, the propagation of the fire front process $X_{t}=[x^{1}_{t},\ldots,x^{N}_{t}]^{\top}\in\mathcal{X}$ is more compactly expressed as: $X_{t}=\xi(X_{t-1},E_{t-1})$, where $E_{t}\sim P_{E}$, with $E_{t}\in[0,2\pi]^{N}\times[0,\infty)^{N}\times[0,\infty)^{N}$, denotes a random realization of $\{\theta(i),w_{s}(i),r_{f}(i)\}_{i=1}^{N}$ drawn from the PDF $P_{E}$, which captures the stochasticity of the fire front propagation acting as a stationary process noise.
@@ II-C Agent Dynamics and Sensing Model
An autonomous mobile agent represented by a point-mass object, evolves inside a bounded planar environment $\mathcal{E}\subset\mathbb{R}^{2}$ according to discrete-time dynamics of the form $y_{t}=f_{a}(y_{t-1},u_{t-1})$ [10], where $y_{t}\in\mathcal{Y}$ is the state of agent at time $t$, and $u_{t}\in\mathcal{U}$ is the control input. In addition, the agent has a finite sensing range for observing its surroundings (i.e., through a camera), which is given by a circular region with radius $R_{a}$ i.e., $O_{t}=\{x\in\mathbb{R}^{2}\mid\|x-y^{p}_{t}\|\leq R_{a}\}$, where $y^{p}_{t}$ is the agent’s position at time $t$.
The agent uses its camera to observe the state of the fire front, i.e., by taking snapshots and determining the location of the fire front from the image snapshots using image processing (i.e., object detection). Due to sensing and image processing imperfections, this process carries a certain degree of inaccuracy, resulting in noisy observations. Specifically, for fire front vertex $i$ with true state $x^{i}_{t}$, the agent observes the measurement $z^{i}_{t}$ inside its sensing range according to: $z^{i}_{t}=h(x^{i}_{t})+w^{i}_{t},$ where $h(\cdot)$ is a function that relates the true states to the received measurements, and $w^{i}_{t}\sim\mathcal{N}(0,\sigma^{2}_{z}I_{2\times 2})$ represents measurement noise. The noise is independent and identically distributed (i.i.d.) according to a zero-mean Gaussian distribution with variance $\sigma^{2}_{z}$, where $I_{2\times 2}$ is the $2\times 2$ identity matrix. Additionally, $w^{i}_{t}$ is independent of the process noise described in the previous section.
@@ III-A Fire Front Recursive State Estimation
where with slight abuse of notation in Eq. (III-A) $x_{t}\in\mathbb{R}^{d_{x}}$ is the state of the system, $z_{t}\in\mathbb{R}^{d_{z}}$ is the received measurement, $p(x_{t}|x_{t-1})$ is the transitional density governed by the stochastic process dynamics, $p(z_{t}|x_{t})$ is the measurement likelihood function, $\mathcal{B}^{-}_{t}(x_{t}|z_{1:t-1})$ is the predictive belief distribution at time $t$, and $\mathcal{B}_{t}(x_{t}|z_{1:t})$ is the posterior belief of $x_{t}$ when all measurements $z_{1:t}=[z_{1},\ldots,z_{t}]$ up to time $t$ have been received. Subsequently, given the recursion in Eq. (III-A) the minimum mean square estimator (MMSE) $\hat{x}^{\text{MMSE}}_{t}$ is given by: $\hat{x}^{\text{MMSE}}_{t}=\int x_{t}\,\mathcal{B}_{t}(x_{t}|z_{1:t})\,dx_{t}$.
However, in our problem, at time $t$ the observation is a point pattern of random cardinality, not a fixed-length vector, hence the standard vector-likelihood underlying Eq. (III-A) is not directly applicable. We therefore replace the likelihood in Eq. (III-A) with the appropriate set-likelihood and apply Bayes’ rule with that form. To achieve this, first observe that the transitional density $p(x_{t}|x_{t-1})$ in Eq. (III-A) becomes $p(X_{t}|X_{t-1})=\int\delta(X_{t}-\xi(X_{t-1},E_{t-1}))\,P_{E}(E_{t-1})\,dE_{t-1}$ as direct consequence of the fire front stochastic dynamics, where $\delta(\cdot)$ is the Dirac delta function. Subsequently, we can compute the predicted belief $\mathcal{B}^{-}_{t}(X_{t}|Z_{1:t-1})$ at time step $t$, assuming the posterior density at the previous time step, $\mathcal{B}_{t-1}(X_{t-1}|Z_{1:t-1})$, is known. The posterior belief $\mathcal{B}_{t}(X_{t}|Z_{1:t})$ can then be obtained by incorporating the measurement set $Z_{t}$ via the correction step shown in Eq. (III-A), provided that an expression for the measurement likelihood function $p(Z_{t}|X_{t},y_{t})$ is available. This process can then be recursively applied to the next time step.
@@
$\frac{1}{m_{t}!}\exp\left(-\sum_{i=1}^{N}\lambda_{i}(x^{i}_{t})\right)\prod_{k=1}^{m_{t}}\left(\,\sum_{i=1}^{N}\gamma^{i}_{t}\bigl(z_{k}|x^{i}_{t},y_{t}\bigr)\right),$ (5)

### v2

@@ I Introduction
The 2025 Southern California wildfires were among the most devastating in the state’s history: the Palisades Fire in Los Angeles and the Eaton Fire in Altadena burned nearly 40,000 acres, destroyed over 16,000 structures, and displaced hundreds of thousands of people [7]. Accurate estimation of wildfire propagation is therefore critical for effective disaster response [18, 16, 10] and informed decision-making [17, 12]. Motivated by this need, this work investigates adaptive monitoring of a stochastic wildfire front using a mobile agent.
The task requires planning the agent’s trajectory over a rolling finite horizon to minimize uncertainty in estimating the fire’s evolution from sensor data. At each step, the agent re-plans based on the current environment, yielding a complex problem that demands an integrated approach to sensing [15, 20, 19], estimation[14, 25, 11] , and control [23, 22, 13]. Existing methods often address only parts of this problem, either by decoupling sensing, estimation , and control, or by simplifying assumptions [26, 30]. The problem considered in this work relates to active sensing [9], information gathering [1], and sensor management [5]. Sensor management approaches typically address stateless sensors without dynamics, focusing on placement or node selection, and are therefore limited in scenarios where sensor states evolve (e.g., drones equipped with onboard sensors). In such dynamic settings, adaptive control strategies are required. For stochastic dynamic processes, many methods adopt myopic control [3], whereas non-myopic schemes are typically greedy [28] or heuristic with sub-optimality guarantees [6]. Approaches that provide optimality guarantees often assume linear dynamics with Gaussian noise and apply the certainty equivalence principle (CEP), thereby reducing the problem to deterministic optimal control [8]. These assumptions, however, do not hold for the problem addressed in this work.
@@
The rest of the paper is organized as follows: Section II formulates the problem, Section III presents the proposed methodology, Section IV provides an evaluation of the approach, and finally, Section V concludes the paper.
@@ II-A Problem Objective
Here, the subscript $t+\tau|t$ denotes predicted quantities at time step $t+\tau$ for $\tau\in\{1,\dots,T\}$ within the planning horizon, based on information available at time step $t$. The cost function $\mathcal{C}_{\tau}$ is a bounded, real-valued function that takes as input: a) the predictive density $\mathcal{B}^{-}_{t+\tau|t}$ i.e., $\mathcal{B}^{-}_{t+\tau|t}(X_{t+\tau|t}|Z_{1:t+\tau-1|t})$, of the fire front state at time $t+\tau|t$, and b) the future i.e., predicted, measurement $Z_{t+\tau|t}(u_{t+\tau-1|t})$, which will be received given that the agent executes the control input $u_{t+\tau-1|t}$ and moves to the predicted state $y_{t+\tau|t}$. It then returns a value representing the uncertainty of the fire front state captured in the resulting (pseudo) posterior belief $\mathcal{B}_{t+\tau|t}$.
An information-rich measurement set is one that reduces the dispersion in the posterior, which is the agent’s objective over the planning horizon. The expectation is taken with respect to the future measurement set $Z_{t}^{1:T}=\{Z_{t+1|t},\ldots,Z_{t+T|t}\}$. Subsequently, the agent executes the first control input in the sequence, i.e., $u_{t|t}$, transitions to its new state $y_{t+1}$, receives the real measurement $Z_{t+1}$, computes the posterior belief $\mathcal{B}_{t+1}(X_{t+1}|Z_{1:t+1})$, and repeats the process described above for time step $t+1$.
@@ II-B Fire Front Propagation Model
where:
$SC(i)$ $=\text{x}^{i}_{s}\sin\left(\theta(i)\right)+\text{y}^{i}_{s}\cos\left(\theta(i)\right),$ (3)
$CS(i)$ $=\text{x}^{i}_{s}\cos\left(\theta(i)\right)-\text{y}^{i}_{s}\sin\left(\theta(i)\right),$
$C_{1}(i)$ $=\alpha_{3}(i)\,\sin\left(\theta(i)\right),$
$C_{2}(i)$ $=\alpha_{3}(i)\,\cos\left(\theta(i)\right),$
and $\text{x}^{i}_{s},\text{y}^{i}_{s}]^{\top}$ are the components of the tangent vector at vertex $i$, providing the local orientation of the fire front at that point.
Environmental conditions, such as fuel type and weather, local to each vertex, affect the forward fire propagation rate and direction. These factors include wind direction and speed, denoted by $\theta$ and $w_{s}$, respectively, as well as the fire spread rate due to fuel type, denoted by $r_{f}$. These parameters are stochastic and can vary throughout the environment. Therefore, each vertex may be affected differently depending on the fire front’s extent and the environmental variability.
Specifically, we denote $\theta(i)\in[0,2\pi]$ as the wind direction affecting vertex $i$, the wind speed at the location of vertex $i$ as $w_{s}(i)\in\mathbb{R}^{+}$, and the fire spread rate as $r_{f}(i)\in\mathbb{R}^{+}$. Here, $\theta(i),w_{s}(i),r_{f}(i)$, $\forall i$, are random realizations of the wind direction, wind speed, and fire spread rate at the location of vertex $i$.
The parameters $\alpha_{1}(i)$, $\alpha_{2}(i)$, and $\alpha_{3}(i)$ denote the shape parameters governing the elliptical fire growth from vertex $i$, representing respectively the lengths of the semi-minor axis, the semi-major axis, and the distance from the ignition point to the center of the ellipse, defined respectively as:
$\alpha_{1}(i)$ $=\frac{r_{f}(i)+\dfrac{r_{f}(i)}{HB(i)}}{2\,LB(i)},$ (4)
$\alpha_{2}(i)$ $=\frac{r_{f}(i)+\dfrac{r_{f}(i)}{HB(i)}}{2},$
$\alpha_{3}(i)$ $=\alpha_{2}(i)-\frac{r_{f}(i)}{HB(i)}.$
where $HB(i)$ is the head-to-back ratio accounting for the difference between the fire’s forward (head) and backward (back) spread from the ignition point, while $LB(i)$ is the length-to-breadth ratio which determines the overall elongation of the fire’s elliptical shape. These are defined as:
$HB(i)$ $=\frac{LB(i)+\bigl(LB(i)^{2}-1\bigr)^{0.5}}{LB(i)-\bigl(LB(i)^{2}-1\bigr)^{0.5}},$ (5)
[…]
For a more in-depth description these parameters we refer the reader to [4]. Subsequently, the propagation of the fire front process $X_{t}=[x^{1}_{t},\ldots,x^{N}_{t}]^{\top}\in\mathcal{X}$ is more compactly expressed as:
$X_{t}=\xi(X_{t-1},E_{t-1}),$ (6)
where $E_{t}\sim P_{E}$, with $E_{t}\in[0,2\pi]^{N}\times[0,\infty)^{N}\times[0,\infty)^{N}$, denotes a random realization of $\{\theta(i),w_{s}(i),r_{f}(i)\}_{i=1}^{N}$ drawn from the PDF $P_{E}$, which captures the stochasticity of the fire front propagation acting as a stationary process noise.
@@ II-C Agent Dynamics and Sensing Model
An autonomous mobile agent represented by a point-mass object, evolves inside a bounded planar environment $\mathcal{E}\subset\mathbb{R}^{2}$ according to discrete-time dynamics of the form [24, 21]:
$y_{t}=f_{a}(y_{t-1},u_{t-1})$ (7)
where $y_{t}\in\mathcal{Y}$ is the state of agent at time $t$, and $u_{t}\in\mathcal{U}$ is the control input. In addition, the agent has a finite sensing range for observing its surroundings (i.e., through a camera), which is given by a circular region with radius $R_{a}$ i.e., $O_{t}=\{x\in\mathbb{R}^{2}\mid\|x-y^{p}_{t}\|\leq R_{a}\}$, where $y^{p}_{t}$ is the agent’s position at time $t$.
The agent uses its camera to observe the state of the fire front, i.e., by taking snapshots and determining the location of the fire front from the image snapshots using image processing (i.e., object detection). Due to sensing and image processing imperfections, this process carries a certain degree of inaccuracy, resulting in noisy observations. Specifically, for fire front vertex $i$ with true state $x^{i}_{t}$, the agent observes the measurement $z^{i}_{t}$ inside its sensing range according to:
$z^{i}_{t}=h(x^{i}_{t})+w^{i}_{t},$ (8)
where $h(\cdot)$ is a function that relates the true states to the received measurements, and $w^{i}_{t}\sim\mathcal{N}(0,\sigma^{2}_{z}I_{2\times 2})$ represents measurement noise. The noise is independent and identically distributed (i.i.d.) according to a zero-mean Gaussian distribution with variance $\sigma^{2}_{z}$, where $I_{2\times 2}$ is the $2\times 2$ identity matrix. Additionally, $w^{i}_{t}$ is independent of the process noise described in the previous section.
@@ III-A Fire Front Recursive State Estimation
where with slight abuse of notation in Eq. (III-A) $x_{t}\in\mathbb{R}^{d_{x}}$ is the state of the system, $z_{t}\in\mathbb{R}^{d_{z}}$ is the received measurement, $p(x_{t}|x_{t-1})$ is the transitional density governed by the stochastic process dynamics, $p(z_{t}|x_{t})$ is the measurement likelihood function, $\mathcal{B}^{-}_{t}(x_{t}|z_{1:t-1})$ is the predictive belief distribution at time $t$, and $\mathcal{B}_{t}(x_{t}|z_{1:t})$ is the posterior belief of $x_{t}$ when all measurements $z_{1:t}=[z_{1},\ldots,z_{t}]$ up to time $t$ have been received. Subsequently, given the recursion in Eq. (III-A) the minimum mean square estimator (MMSE) $\hat{x}^{\text{MMSE}}_{t}$ is given by:
$\hat{x}^{\text{MMSE}}_{t}=\int x_{t}\,\mathcal{B}_{t}(x_{t}|z_{1:t})\,dx_{t}.$ (11)
However, in our problem, at time $t$ the observation is a point pattern of random cardinality, not a fixed-length vector, hence the standard vector-likelihood underlying Eq. (III-A) is not directly applicable. We therefore replace the likelihood in Eq. (III-A) with the appropriate set-likelihood and apply Bayes’ rule with that form. To achieve this, first observe that the transitional density $p(x_{t}|x_{t-1})$ in Eq. (III-A) becomes:
$p(X_{t}|X_{t-1})=\int\delta(X_{t}-\xi(X_{t-1},E_{t-1}))\,P_{E}(E_{t-1})\,dE_{t-1},$
as direct consequence of the fire front stochastic dynamics, where $\delta(\cdot)$ is the Dirac delta function. Subsequently, we can compute the predicted belief $\mathcal{B}^{-}_{t}(X_{t}|Z_{1:t-1})$ at time step $t$, assuming the posterior density at the previous time step, $\mathcal{B}_{t-1}(X_{t-1}|Z_{1:t-1})$, is known. The posterior belief $\mathcal{B}_{t}(X_{t}|Z_{1:t})$ can then be obtained by incorporating the measurement set $Z_{t}$ via the correction step shown in Eq. (III-A), provided that an expression for the measurement likelihood function $p(Z_{t}|X_{t},y_{t})$ is available. This process can then be recursively applied to the next time step.
@@
$\frac{1}{m_{t}!}\exp\left(-\sum_{i=1}^{N}\lambda^{i}_{t}(x^{i}_{t})\right)\prod_{k=1}^{m_{t}}\left(\,\sum_{i=1}^{N}\gamma^{i}_{t}\bigl(z_{k}|x^{i}_{t},y_{t}\bigr)\right),$ (12)
@@
Proof:
Due to the independence of noise realizations in the measurement process, the measurements are conditionally independent given the fire front state $X_{t}$. As a result, the Poisson processes defined by Eq. (9) are themselves independent, which implies that the combined set of all measurements generated by all processes forms a superposition of Poisson point processes, with total intensity $\Gamma_{t}(z|X_{t},y_{t})=\sum_{i=1}^{N}\gamma^{i}_{t}(z|x^{i}_{t},y_{t})$, and probability density function (PDF) $p(z|X_{t})=\Gamma_{t}(z|X_{t},y_{t})\left(\int_{O_{t}}\Gamma_{t}(z|X_{t},y_{t})\,dz\right)^{-1}=\Gamma_{t}(z|X_{t},y_{t})\Lambda_{t}^{-1}$. Subsequently, the joint likelihood $p(Z_{t}|X_{t},y_{t})$ of receiving $m_{t}$ measurements at time $t$ can be decomposed as $p(Z_{t}|X_{t},y_{t})=p_{m}(m_{t})\,p(z^{1}_{t},\ldots,z^{m_{t}}_{t}|m_{t},X_{t},y_{t})$, where $p_{m}(m_{t}|y_{t})$ is the probability of receiving $m_{t}$ observations inside the sensing range, and $p(z^{1}_{t},\ldots,z^{m_{t}}_{t}|m_{t},X_{t},y_{t})$ is the conditional joint likelihood function. The term $p_{m}(m_{t},|y_{t})$ follows a Poisson distribution with parameter $\Lambda_{t}$, i.e., $p_{m}(m_{t}|y_{t})={m_{t}!}^{-1}\exp(-\Lambda_{t})\Lambda_{t}^{m_{t}}$. In addition, the joint likelihood decomposes as $p(z^{1}_{t},\ldots,z^{m_{t}}_{t}|m_{t},X_{t},y_{t})=\prod_{k=1}^{m_{t}}p(z_{k}|X_{t},y_{t})$, and thus $\prod_{k=1}^{m_{t}}p(z_{k}|X_{t},y_{t})=\prod_{k=1}^{m_{t}}\Gamma_{t}(z_{k}|X_{t},y_{t})\Lambda_{t}^{-1}$. Since each vertex only contributes locally to its own measurement process, it follows that $\prod_{k=1}^{m_{t}}\Gamma_{t}(z_{k}|X_{t},y_{t})\Lambda_{t}^{-1}=\prod_{k=1}^{m_{t}}\left(\sum_{i=1}^{N}\gamma^{i}_{t}(z_{k}|x^{i}_{t},y_{t})\Lambda_{t}^{-1}\right)$. Since $\int_{O_{t}}\Gamma_{t}(z|X_{t},y_{t})dz$ = $\sum_{i=1}^{N}\int_{O_{t}}\gamma^{i}_{t}(z|x^{i}_{t},y_{t})dz$ = $\sum_{i=1}^{N}\lambda^{i}_{t}(x^{i}_{t})$, the result in Eq. (12) follows directly. ∎